OpenStudy (anonymous):

solve for x: log(1-x) - log(8+x) = 1 I got this so far: (1-x/8+x) = 10

OpenStudy (anonymous):

multiply both sides by (8+x) and post what you get.

OpenStudy (anonymous):

ok. (8/16+x^2) = 80+10x

OpenStudy (anonymous):

thats wrong i know it

OpenStudy (anonymous):

yup, it is catastrophically wrong.

OpenStudy (anonymous):

ok im retarded maybe its the caffeine buz

OpenStudy (anonymous):

please explain why u have to multiply by 8+x on both. why not 1-x on both?

OpenStudy (anonymous):

always be clear in what you are writing. use brackets to make your meaning clear. it is (1-x)/(8+x) = 10 multiply by (8+x) on both sides (1-x)*(8+x)/(8+x) = 80+10x the (8+x) cancels out in the numerator and denomiator on the left hand side, leaving 1-x = 80+10x. I trust you can take it from there?

OpenStudy (anonymous):

Yes thanx

OpenStudy (anonymous):

x=79/11

OpenStudy (anonymous):

you multiply by 8+x since you want to cancel it out from the denominator.

OpenStudy (anonymous):

- negative

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

dude i just noticed, im so dumb. haha